OpenStudy (anonymous):

4.4 x 10 to the 7 power, my answer is 440,000,000 is this correct?

OpenStudy (anonymous):

4.4 x (10 to the 7) = 44000000

OpenStudy (anonymous):

i dont think she is correct

OpenStudy (anonymous):

yea i just looked closely

OpenStudy (anonymous):

44,000,000

OpenStudy (unklerhaukus):

\[4.4\times10^7=44\times10^6=44,000,000\checkmark\]

OpenStudy (anonymous):

I have to move the decimal to the left by one and every thing else is okay?

OpenStudy (anonymous):

Thank you, everyone!!!

OpenStudy (unklerhaukus):

yeah, if you want to do it in steps \[4.4\times10^7\\=44\times10^6\\=440 \times10^5\\=4400 \times10^4\\=44000 \times10^3\\=440000 \times10^2\\=4400000 \times10\\=44000000\\\hspace{-.4em}=44,000,000\]

OpenStudy (anonymous):

Thank you very helpful....